OpenStudy (anonymous):

Hi guys I wanted to know which one of these are seperable. and whats the equation and i want to do the rest please help out. 1) dy/dx = (xy+y)/4x-3xy) 2) dy/dx = y^2 +2x +3y 3) dy/dx = ye^(sinx + cosy)

OpenStudy (anonymous):

If it's separable, it may be rearranged into this form \[f(y)dy = g(x)dx\]

OpenStudy (anonymous):

So nothing with x multiplies dy, and nothing with y multiplies dx.

OpenStudy (anonymous):

yeah i know 3 is seperable i got that but i got confused by 1 and 2 the most

OpenStudy (sirm3d):

1 is separable. \[\frac{xy+y}{4x-3xy}=\frac{y(x+1)}{x(4-3y)}=\frac{y}{4-3y}\cdot\frac{x+1}{x}\]